OpenStudy (anonymous):

mix 2 liters of 30c water with 3 litters of 50c water and you have 5 litters of water at what degrees?

OpenStudy (whpalmer4):

Don't you just convert to absolute temperature (\(^\circ\text{K}\)) and take a weighted average?

OpenStudy (anonymous):

no its still in C

OpenStudy (whpalmer4):

I understand that your temperatures are currently in degrees C. My point is that to compute the final temperature, you convert both temperatures to degrees K, then take a weighted average. That will give you the temperature (in degrees K) of the mixture.

OpenStudy (whpalmer4):

Do you understand what I mean by "take a weighted average"?

OpenStudy (anonymous):

yes but the answer is in degrees

OpenStudy (whpalmer4):

And if you do as I suggest, what do you get?

OpenStudy (anonymous):

i get 43

OpenStudy (anonymous):

?

OpenStudy (whpalmer4):

First, what do you get when you convert \(30 ^\circ\text{C}\) to \(^\circ\text{K}\)?

OpenStudy (anonymous):

lets do it together

OpenStudy (whpalmer4):

You already got an answer, I want to see what you did.

OpenStudy (anonymous):

may we just do it together?

OpenStudy (whpalmer4):

What do you mean by that? I do the problem, and show you what I did?

OpenStudy (anonymous):

Yeah

OpenStudy (whpalmer4):

Okay, sure, I'll do that, but I'll not help you again. To convert \(^\circ\text{C}\) to \(^\circ\text{K}\), add \(273.15^\circ\text{K}\) and change the unit from \(^\circ\text{C}\) to \(^\circ\text{K}\) \[30^\circ\text{C} = (30+273.15) ^\circ\text{K} = 303.15^\circ\text{K}\]\[50^\circ\text{C} = (50+273.15) ^\circ\text{K} = 323.15^\circ\text{K}\] Compute the weighted average: \[\frac{2\text{ liters}}{5\text{ liters}}*303.15^\circ\text{K} + \frac{3\text{ liters}}{5\text{ liters}}*323.15^\circ\text{K} = \] Convert your answer back to \(^\circ\text{C}\) by subtracting \(273.15^\circ\text{K}\) and changing the unit to \(^\circ\text{C}\).